# Deep Phase Coded Image Prior

### 05 Apr 2024 v1

@@ 3 Deep Phase Coded Image Prior
Figure 3: A qualitative comparison of image reconstruction. Our method produced a much more accurate all-in-focus images, The existing baseline produced over-smooth all-in-focus images compared to our method.
4 Experimental Results
4.1 Simulation Results
To evaluate the performance of our method quantitatively both in terms of depth estimation and all-in-focus image reconstruction, we first perform an evaluation on a simulated dataset, namely, ‘TAUAgent’ dataset [^1] https://www.cs.toronto.edu/ harel/TAUAgent/home.html [4]. The Agent dataset consists of a total of 530 images from five synthetic scenes created using the ‘Blender’ computer graphics software (named ”City”, ”Headbutt”, ”Sitting”, ”WallSlam”, ”WuManchu”). Each scene consists of a ground-truth low-noise all-in-focus image, along with its corresponding pixel-wise accurate depth map. Such data enables an exact depth-dependent imaging simulation, with the corresponding depth of focus (DOF) effects. In contrast to prior works [7, 4, 3], we do not use the ‘ground-truth’ data from the Agent dataset in our optimization. We only use the acquired coded images produced by the camera simulation. The all-in-focus GT images and the GT depth maps are only used to evaluate the performance of our method. For the all-in-focus evaluation, we used PSNR and SSIM, while for the depth estimation, we used average error in meters (as done in [3, 4]). To make a fair comparison we simulate the imaging with the same optical parameters as done in [7, 3, 4].
Note that a crucial disadvantage of previous works is the high dependency on a large and high-quality dataset for successful learning. High-quality datasets like the ‘agent’ dataset are usually small and cumbersome to collect, thus, making the learning much more challenging. In contrast to them, our Deep Phased implicit representation reduces this constraint since we explicitly incorporate the DCM into the learning scheme. Thus, we can extract both the all-in-focus RGB image and the pixel-level depth map from a single captured image.
Figure 4: Blind image deblurring as a function of the kernel size. The graph shows the average performance on a subset of samples to choose the best kernel size for Nueral deblurring on the dataset. We can see that as the kernel size drops the performance improves, up to a certain level where the loss of deblurring signal is too large.
We took a subset of 110 images (namely, “agent subset”) from the scenes mentioned above to evaluate the performance on several tasks namely passive depth estimation, image reconstruction, and image deblurring. Lastly, we also show the effect of the implicit generator on the performance of the mentioned tasks.
4.1.1 Passive Depth Estimation
To assess our method’s performance in passive depth estimation we compared ourselves to the Monocular depth estimation network for phase coded imaging from [4] which built upon the same phase-coded imaging system. We also followed the same optical parameters to produce the same acquired images. The evaluation was done on the relevant part of the ”agent subset” (”WuManchu” scene) that was not used for training [4]. Table I shows a quantitative comparison in terms of RMSE error in meters. Visual results can be found in Figures 2 ,7. Overall, our method outperformed mono network by two orders of magnitude and effectively managed to produce an accurate pixel-level depth map solely from a given captured image.
TABLE II: A quantitative comparison of for image reconstruction (PSNR [dB] / SSIM). Our method outperforms the existing supervised baseline significantly. The PSNR average improvement of our method is $\sim 5dB$.
Supervised 0-shot
Dataset Baseline [7] DPCIP
City 24.01/0.65 29.73/0.91
Headbutt 28.54/0.92 31.28/0.94
Sitting 26.12/0.81 30.04/0.91
WallSlam 26.94/0.87 31.45/0.93
WuManchu 23.67/0.71 30.86/0.93
4.1.2 Extended Depth Of Field (EDOF)
To assess our method performance in the task of extended depth of field (EDOF) we compared ourselves to [7] which built upon the same phase-coded imaging system, we also follow the same optical parameters to produce the same acquired images. In this case, the original method was trained on a different synthetic dataset, so we could test on all the images in our subset. Table II shows a quantitative comparison in terms of PSNR and SSIM. Visual results can be found in Figure 3. Overall, our method outperformed the baseline by 5 dB in PSNR and 0.13 in SSIM on average and showed an improvement in terms of image reconstruction metrics, remarkably, relying only on a single given phase-coded image. Our method outperformed the previous supervised method [7] both in terms of PSNR [dB] and SSIM with the same imaging system. Table II shows a quantitative comparison using the agent subset. Qualitative results can be found in Figures 3, 8.
TABLE III: PSF Kernel energy spread analysis. Compared to the full signal, the Blue kernel is the most centered, while the Red kernel is the most spread out. We can see that reducing the kernel by half nearly effect the energy, but as we keep reducing the kernel size The loss of energy becomes severe.
Kernel size R G B
71x71 100% 100% 100%
31x31 94.72% 96.69% 98.49%
23x23 87.31% 94.30% 97.32%
19x19 77.20% 89.10% 96.25%
15x15 56.24% 78.81% 94.28%
GT Input Neural-Deblurring DPCIP
@@ 4.1.2 Extended Depth Of Field (EDOF)
Figure 5: Image deblurring. Our method outperforms Neural Deblurring. Even though we adapt Neural deblurring to be able to handle our deblurring subset, our method still outperforms by a significant margin, making it more usable for real-world applications where the PSF kernels can be large.
4.1.3 Deblurring
We extend our experiments to solve blind image deblurring. The problem of image deblurring can be considered as a specific case of reconstructing an all-in-focus image from defocus where the full captured image is approximately at the same distance. We can simulate such a setup by taking images that the whole scene is above the physical distance of the camera, effectively causing the phase-coded imaging to collapse into a uniform convolution with a single PSF kernel.
For that purpose we also compared our method with Neural Deblurring [13], a similar approach that uses DIP as a generator and a fully-connected network (FCN) to recover the sharp image and the blurring kernel respectively. For the purpose of deblurring we took from the ‘agent’ dataset 13 examples from 3 scenes. Each example has a constant depth map as described before.
However, Neural Deblurring fatally fails in scenarios where the deblurring kernel is large, and also when it’s channel dependent. Since comparing to Neural deblurring ’out of the box‘ is not really relevant, We adapt Neural Deblurring to work ith our setup (e.g. phase mask kernel sizes), First we let it to predict each channel separately to match the models’ capacity to our problem. In addition, to overcome the kernel size limitation, we effectively reduced the predicted kernel from the original size of 71x71 to 19x19. The reduction in size can be done since the pre-defined PSF kernels are centered.
In Table III we show a quantitative analysis of the signal coverage of reduced kernels from the full PSF kernels. In the trade-off between the energy loss of the PSF kernels and the ability of Neural Deblurring to converage, We chose kernel size of 19x19 to predict on each channel independently and average the results to compare vs ur method. Figure 4 demonstrates the deblurring improvement as a function of the reduction in kernel size.
To conclude our blind image deblurring experiments, Table V summarize our results quantitatively and Figures 5 and 9 demonstrate the results qualitatively for the deblurring case. Overall we can see that our method manages to outperform in blind image deblurring by effectively recovering a single valued depth map.
@@ 4.1.4 Implicit generator effect
TABLE IV: Generator effect. We evaluated three types of generators in our method. DIP outperformed other generators in image reconstruction and was comparable to PIP in depth estimation. Both DIP and PIP outperformed SIREN in both tasks.
@@ 4.2 Real-World Results
We further extend our simulation results to real-world examples using a dedicated camera as done in [3, 7, 4]. For each image crop we compare ourselves to the relevant baseline in image reconstruction and passive depth estimation. Figures 6 and 10 summarizes our results.
Interestingly, DIP which dominates in the simulation setup for all datasets and both tasks, is not working as well in the real-world setup. A possible hypothesis that can explain the results is that DIP is not as good as PIP at disentangling depth from texture, resulting in incorrect depth values and as a result, incorrect reconstruction.
In the simulated data there is less texture since we have no noise, thus the image is much more like a smooth depth map. To overcome the noise effect, we applied TV regularization to the estimated depth map to help the model produce a smoother depth map. The smoothness regularization helped achieve better depth maps but it is still not as good as PIP with the same regularization. When using PIP as our generator, we have more control over the smoothness by specifying an initialization scheme for the frequencies and limiting their bandwidth.
Overall, setting up the original settings (i.e. log-scale frequencies initializations and limiting the range as done in [22]) with TV regularization produced the best results.
Blur img Baseline DPCIP (DIP + TV Reg) DPCIP (PIP + TV Reg)
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Real world examples. We can see that using PIP as the generator generates the most accurate depth maps and also the best all-in-focus images. Using DIP as a generator tends to produce too detailed depth maps, but with good all-in-focus images, this means that PIP as a generator disentangle better between the depth map estimation and the all-in-focus restoration. In all cases, previous baselines produced worse outputs compare to our method. depth maps are over-smoothed and all-in-focus images are a bit noiser than ours.
@@ 5 Conclusion
In this work, we have proposed a self-supervised method for recovering an absolute depth map and an all-in-focus image using a phase-coded imaging system. As opposed to previous methods that rely on datasets for recovering the depth map or the all-in-focus image, we do it solely on the given captured image and the knowledge of the optical setup by leveraging the image prior of implicit models.
Our method reduces the need for using training datasets since collecting such datasets can be hard and requires external graphics rendering software. Providing a solution for real-world setups and users who seek to use an existing imaging system without the need for any additional information or data other than optical parameters. In addition, due to the high quality of restoration, our method can provide ”pseudo GT” to be used as supervision for training a neural network to restore an all-in-focus image from the blurred acquired image. This way we can enjoy both high quality of reconstruction and fast inference time.
TABLE V: Image deblurring. Our method outperforms previous blind image deblurring works (even after the adaptation) both in PSNR [dB] and SSIM.
Baseline [13] DPCIP
Deblurring 19.6/0.67 26.38/0.834

### 09 Mar 2025 v2

@@ 4.1.2 Extended Depth Of Field (EDOF)
Figure 4: A qualitative comparison of image reconstruction. Our method produced a much more accurate all-in-focus images. The existing EDOF baseline [7] produced over-smooth all-in-focus images compared to our method.
4.1.3 Deblurring
We extend our experiments to solve blind image deblurring. For image deblurring, we aim to reconstruct a sharp image from a blurred one assuming we have a single blur kernel. Image deblurring can be thought of as a reduction of the defocus problem as in defocus we have a depth-dependent blur kernel. Since our imaging system has an effective range, everything outside the effective range will be clipped to the range’s limits. Thus we can simulate a deblurring setup from our defocus problem and the system’s physical limitations. To simulate a deblurring experiments, we chose images where ground-truth depth values are all outside the imaging system’s effective range, values outside the effective range will result a clip of the $\psi$ values, hence we will have acquired images with a single PSF blur kernel- which is a exactly a deblurring scenario.
We compare our method with Neural Deblurring [13], a single image deblurring approach that shares several key components with our method, such as using DIP as a generator and a learned module (fully-connected network) to recover the sharp image and the blurring kernel respectively. As mentioned, to mimic a deblurring setup we took from the ‘agent’ dataset 13 examples from 3 scenes. Each example has a constant depth map ($\psi=-4$). We ran DPCIP and Neural Deblurring on all images and compared PSNR and SSIM for image reconstruction. However, Neural Deblurring fatally fails in scenarios where the deblurring kernel is large (the PSF kernels in the DCM are 71x71). In addition, our blur kernels are channel-dependent. Since comparing to Neural deblurring ‘out of the box’ is not practical, we modify Neural Deblurring to work with our setup (e.g. phase mask kernel sizes). First, we let it predict each channel separately to match the models’ capacity to our problem. In addition, to overcome the kernel size limitation, we effectively reduced the predicted kernel from the original size of 71x71 to 19x19. The size reduction can be done since the pre-defined PSF kernels are centered (See more details in the supplementary material).
To conclude our blind image deblurring experiments, Table III summarizes our results quantitatively and Figure 5 demonstrates the results qualitatively for the deblurring case (more examples can be found in the supplementary materials). Overall we can see that our method manages to outperform in blind image deblurring by effectively recovering a single-valued depth map.
0-shot 0-shot
GT Input Neural-Deblurring[13] DPCIP
@@ 4.1.3 Deblurring
Figure 5: Image deblurring. Our method outperforms Neural Deblurring [13] Even though we adapt Neural Deblurring to our deblurring subset, our method still outperforms it by a significant margin, making our approach more usable for real-world applications where the PSF kernels can be large.
TABLE III: Image deblurring. Our method outperforms previous blind image deblurring works (even after adaptation) both in PSNR [dB] and SSIM.
Neural Deblurring [13] DPCIP
Deblurring 19.6/0.67 26.38/0.834
4.2 Ablation studies
@@ 4.2.1 Implicit generator effect
TABLE IV: Generator effect. We evaluated three types of generators in our method. DIP outperformed other generators in image reconstruction and was comparable to PIP for depth estimation. Both DIP and PIP outperformed SIREN in both tasks.
@@
4.2.2 SSIM Loss effect
We justify the switch from L2 loss to SSIM during the optimization using the following ablation study. We collected 15 images from 3 scenes to be used as our test set. We ran the same optimizations on all images, with and without the switch to SSIM after 500 iterations. Table V concludes the effect. We can see that both for image reconstruction and depth estimation switching from L2 to SSIM greatly improved our overall results.
TABLE V: SSIM Loss effect. The SSIM loss improved the results for all image reconstruction metrics (PSNR and SSIM) and also for the depth estimation.
PSNR [dB] SSIM Depth error [m]
L2 25.97 0.813 0.008
L2+SSIM 30.20 0.920 0.003
4.2.3 Mismatch in the optical system
We analyze the robustness of our method with respect to a mismatch in the optical parameters. We took the same 15-image test set used throughout the ablation study. We ran our method several times on the same test set of images, where each time our method had slightly different PSF kernels. The mismatch was done by changing the phase mask, each time by a different factor ($1\%-10\%$). We can see in Table VI that with small mismatches in the PSF, the generations of our method are still fairly robust.
TABLE VI: Analysis of mismatch in the optical system (mismatch in the phase of the mask) Our model is fairly robust to small errors in the optical model.
Phase change $0\%$ $1\%$ $2\%$ $5\%$ $10\%$
PSNR[dB] / SSIM 30.15/0.916 29.57/0.911 27.90/0.88 25.90/0.83 24.41/0.77
@@ 4.3 Real-World Results
We further extend our simulation results to real-world examples using a dedicated camera as done in [3, 7, 4]. For each image crop, we compare ourselves to the relevant baseline in both image reconstruction and passive depth estimation. Figure 6 provide a qualitative comparison of the results.
Since real-world images suffer from much higher noise levels we add TV regularization to help DPCIP to better estimate the depth maps. For each image we ran DPCIP twice, with DIP and PIP [22] as generators. PIP has the ability to control over the initialization scheme of the input code. From the visual results in Figure 6 we can see that PIP manages to better disentangle between the depth maps and the image content, thus leading to better results both in terms of image reconstruction and depth maps.
We further extend our experiments with our dataset to [44], which train a supervised network to extract depth from defocus cues. The results on our dataset are worse than our method or [4]. An explanation for the degradation in performance is needed for a large dataset. In the original paper, the training dataset was ×$\times 10$ more than our dataset, which strengthens DPCIP’s advantages as a zero-shot method that does not rely on a training dataset at all. The supplementary material contains visual examples. In the visual results we can see that regarding depth estimation, the supervised baseline [4] produced smoother depth maps, but can also produce regions with incorrect depths, as opposed to our zero-shot method which has noisier depths but significantly smaller areas with incorrect depths. Compared to the image reconstruction supervised baseline, our zero-shot method produces sharp images that are less noisy. See Figure 7 for zoom-in visualization.
Overall, even though supervised methods are considered as upper-bound, our zero-shot method produced comparable or even better results without the need for any dataset for training which can be considered a significant advantage when it comes to real-world scenarios where the limited amount of existing datasets might not suffice for training a suitable supervised method. More viusal examples can be found at the supplementary materials.
Supervised 0-shot 0-shot
Blur input image EDOF[7] / Mono from [4] DPCIP (DIP) DPCIP (PIP)
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Supervised 0-shot 0-shot
Blur input image EDOF[7] / Mono from [4] DPCIP (DIP) DPCIP (PIP)
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 6: Real world examples. We can see that using PIP as the generator generates the most accurate depth maps and also the best all-in-focus images. Using DIP as a generator tends to produce too detailed depth maps, but with good all-in-focus images. This means that PIP as a generator disentangles better between depth map estimation and all-in-focus restoration. In all cases, our zero-shot method is comparable to or better than the supervised baselines. For depth estimation, the mono network from [4] seems to produce large areas with incorrect depths while our method manages to be more consistent. In the all-in-focus task, the EDOF network [7] produced a bit noisier results than ours; see Figure 7 for zoom-in visualizations.
Supervised 0-shot 0-shot
Blur img EDOF[7] / Mono from [4] DPCIP (DIP) DPCIP (PIP)
@@
Figure 7: Zoom-in reconstruction results We can see that DPCIP produced smoother results, but kept the details sharp as opposed to EDOF competitor that generated noisier reconstruction results.
@@ 5 Conclusion
In this work, we have proposed a self-supervised method for recovering an absolute depth map and an all-in-focus image using a phase-coded imaging system. As opposed to previous methods that rely on datasets for recovering the depth map or the all-in-focus image, we do it solely on the given captured image and the knowledge of the optical setup by leveraging implicit models as the image prior.
Our method reduces the need for using training datasets since collecting such datasets can be hard and requires external graphics rendering software. In addition, the current datasets can suffer from generalization and domain gap issues. Providing a solution for real-world setups and users who seek to use an existing imaging system without the need for any additional information or data other than optical parameters. In addition, due to the high quality of restoration, our method can provide “pseudo GT” to be used as supervision for training a neural network to restore an all-in-focus image from the blurred acquired image. This way we can enjoy both high quality of reconstruction and fast inference time.
@@ Acknowledgments
We would like to thank Shay Elmalem for his useful comments and advice in addition to providing us the simulation code we based our work on.